# Splat-Nav: Safe Real-Time Robot Navigation in Gaussian Splatting Maps

### v1

@@ A Planning Computation Times
Since Splat-Plan requires the use of A* to initialize a path to generate the safety polytope corridors, it will necessarily be slower than these A* baselines. However, we note that in the Full Trajectory regime, the dominating cost is the calculation of the B-spline due to the number of polytopes in the constraints, with many facets per polytope. This QP was solved using CVXPY with the Clarabel solver. We note that the computation time associated with the B-spline computation procedure can be reduced using embedded solvers or faster programming languages.
TABLE IV: Comparison of Full Trajectory planning algorithm computation times.
@@
Since the polytope and A* computation procedures are relatively fast, while computation of the B-spline is relatively slow, we can compute the B-spline for some small time step in the future, in a receding-horizon manner. In our experiments, this amounted to fitting a B-spline through the current and next polytope. Since the Full Trajectory method finds a connected series of polytopes and is therefore always able to fit a B-spline connecting the start position to goal position, the Local Trajectory method can also still guarantee safety and feasibility since the entire constraint set is separable based on time, at the expense of the optimality of the solution.
@@
TABLE V: Comparison of Local Trajectory planning algorithm computation times.
[…]
B Pose Estimation: Comparison to Gradient-Based Pose Estimators
We compare the performance of the pose estimation algorithms in this work to a gradient-based pose estimator amenable to Gaussian Splatting, termed iNeRF-GS, similar to iNeRF [52], a canonical pose estimator amenable to NeRFs. We note that publicly-available codebases for Gaussian Splatting do not provide analytical gradients required for gradient-based pose estimation. In addition, Gaussian Splatting does not support automatic differentiation. As a result, we leverage finite differences to estimate the gradient of the photometric loss function utilized in the pose estimator (please refer to [52] for additional details on gradient-based pose estimation in radiance fields). We note that our implementation might not be particularly fast, given the approach utilized in estimating the gradients. We provide the summary statistics of the error in the pose estimates computed by each algorithm, in addition to the computation time on a trial with $100$ frames in the Statues scene in Table VI. We note that all methods had a perfect success rate in this problem. The iNeRF-GS algorithm achieves the lowest accuracy and requires the greatest computation time, unlike Colored-ICP, Splat-Loc-SIFT, and Splat-Loc-Glue, which achieve much-higher accuracy with a rotation error less than a degree and a translation error less than 1515cm. iNeRF-GS requires a computation time of about $36.15$ secs per frame, which is about two orders of magnitude slower than the next-slowest method ICP, which requires a computation time of about $110$ milliseconds. Colored-ICP, Splat-Loc-SIFT, and Splat-Loc-Glue require less than $100$ milliseconds of computation time. Compared to all methods, Splat-Loc-Glue yields pose estimates with the lowest mean rotation and translation error, less than $0.06$ deg. and 44mm, respectively, and achieves the fastest mean computation time, less than $42$ milliseconds.
TABLE VI: Comparison of pose estimation algorithms in the Statues scene with $\delta_{R}=20^{\circ}$ and $\delta_{t}=0.1\,\mathrm{m}$. All methods were successful in this problem.
Algorithm Rotation (deg.) Translation Error (m) Computation Time (secs.)
ICP [4] $7.306\mathrm{e}^{1}\pm 4.584\mathrm{e}^{1}$ $1.285\mathrm{e}^{0}\pm 7.496\mathrm{e}^{-1}$ $1.064\mathrm{e}^{-1}\pm 1.918\mathrm{e}^{-2}$
Colored-ICP [34] $8.272\mathrm{e}^{-1}\pm 3.718\mathrm{e}^{-1}$ $1.305\mathrm{e}^{-2}\pm 5.973\mathrm{e}^{-3}$ $4.328\mathrm{e}^{-2}\pm 9.697\mathrm{e}^{-3}$
Splat-Loc-SIFT (ours) $8.590\mathrm{e}^{-2}\pm 6.050\mathrm{e}^{-2}$ $5.590\mathrm{e}^{-3}\pm 7.441\mathrm{e}^{-3}$ $6.327\mathrm{e}^{-2}\pm 2.384\mathrm{e}^{-3}$
Splat-Loc-Glue (ours) $\bm{5.036\mathrm{e}^{-2}\pm 3.088\mathrm{e}^{-2}}$ $\bm{3.298\mathrm{e}^{-3}\pm 2.692\mathrm{e}^{-3}}$ $\bm{4.119\mathrm{e}^{-2}\pm 7.319\mathrm{e}^{-2}}$
iNeRF-GS $1.217\mathrm{e}^{2}\pm 3.371\mathrm{e}^{1}$ $2.444\mathrm{e}^{0}\pm 9.116\mathrm{e}^{-1}$ $3.615\mathrm{e}^{1}\pm 5.436\mathrm{e}^{0}$
C Recursive Pose Estimation
Table VII presents the results in the Flightroom scene with $\delta_{R}=50^{\circ}$ and $\delta_{t}=0.2\,\mathrm{m}$. We note that the Gaussian Splatting representation of the Flightroom scene was of much lower quality, which negatively impacted the performance of many of the pose estimation algorithms. As a result, we attempt to mitigate its impact by ensuring that the robot captures informative images during the duration of the problem. ICP and Colored-ICP always succeed but yield pose estimates of notably low accuracy, with mean rotation errors of about 120 deg and 125 deg and translation errors of about 11.59 m and 0.73 m, respectively. In contrast, Splat-Loc-SIFT fails to produce an estimate in all trials. However, Splat-Loc-Glue achieves a perfect success rate, yields high-accuracy pose estimates with a mean rotation error of about 0.3 deg and a mean translation error less than 5 mm, and takes an average of 37.74 ms per frame. Figure 9 shows the errors in the pose estimates for a single trial.
@@ C Recursive Pose Estimation
TABLE VII: Comparison of pose estimation algorithms in the Flightroom scene with $\delta_{R}=50^{\circ}$ and $\delta_{t}=0.2\,\mathrm{m}$
@@
Algorithm Rotation (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$) ICP [4] $212\mathrm{e}^{2}\pm 2.957\mathrm{e}^{1} 1.1$$159\mathrm{e}^{1}\pm 2.601\mathrm{e}^{1}$ $1.451\mathrm{e}^{-1}\pm 1.471\mathrm{e}^{-1}$ $100$ Colored-ICP [34] $1.253\mathrm{e}^{2}\pm 3.188\mathrm{e}^{1}$ $7.354\mathrm{e}^{-1}\pm 2.540\mathrm{e}^{-1}$ $4.847\mathrm{e}^{-1}\pm 5.472\mathrm{e}^{-2}$ $100$ Splat-Loc-SIFT (ours) - - - 0 Splat-Loc-Glue (ours) $\bm{2.539\mathrm{e}^{-1}\pm 1.543\mathrm{e}^{-1}}$ $\bm{3.810\mathrm{e}^{-3}\pm 2.618\mathrm{e}^{-3}}$ $\bm{3.774\mathrm{e}^{-2}\pm 8.522\mathrm{e}^{-4}}$ $100$
Figure 9: Flightroom Scene with $\delta_{R}=50^{\circ}$ and $\delta_{t}=0.2\,\mathrm{m}$—Pose error and computation time per frame achieved by each pose estimation algorithm. Splat-Loc-SIFT was not successful on any trial and is thus omitted in the figure. While ICP and Colored-ICP fail to provide accurate estimates of the robot’s pose, Splat-Loc-Glue computes high-accuracy pose estimates.
Here, we examine the performance of each pose estimation algorithm in a synthetic scene Stonehenge, beginning with initial rotation and translation errors of $20$ deg. and 0.10.1m. Table VIII shows the performance of each algorithm across the ten trials. Similar to the results observed with the real-world scene, Splat-Loc-SIFT and Splat-Loc-Glue yield high-accuracy pose estimates with average rotation and translation errors less than $0.5$ deg and 55mm respectively, unlike the ICP and Colored-ICP which produce low-accuracy estimates. Moreover, ICP and Colored-ICP seem to diverge in some runs as the estimation procedure proceeds over time, depicted in Figure 10. In contrast, Splat-Loc-Glue achieves a perfect success rate, in contrast with Splat-Loc-SIFT, which achieves a $90\%$ success rate, in addition to attaining the fastest computation time.
TABLE VIII: Comparison of pose estimation algorithms in the Stonehenge scene with $\delta_{R}=20^{\circ}$ and $\delta_{t}=0.1\,\mathrm{m}$.
Algorithm Rotation (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$) ICP [4] $9.702\mathrm{e}^{1}\pm 3.006\mathrm{e}^{1}$ $1.085\mathrm{e}^{0}\pm 7.958\mathrm{e}^{-1}$ $3.604\mathrm{e}^{-1}\pm 5.036\mathrm{e}^{-2}$ $100$ Colored-ICP [34] $7.185\mathrm{e}^{1}\pm 3.224\mathrm{e}^{1}$ $5.175\mathrm{e}^{-1}\pm 2.162\mathrm{e}^{-1}$ $3.678\mathrm{e}^{-1}\pm 6.378\mathrm{e}^{-2}$ $100$ Splat-Loc-SIFT (ours) $1.962\mathrm{e}^{-1}\pm 2.690\mathrm{e}^{-2}$ $3.033\mathrm{e}^{-3}\pm 4.268\mathrm{e}^{-4}$ $1.399\mathrm{e}^{-1}\pm 5.727\mathrm{e}^{-3}$ $90$ Splat-Loc-Glue (ours) $\bm{2.114\mathrm{e}^{-1}\pm 2.265\mathrm{e}^{-1}}$ $\bm{2.962\mathrm{e}^{-3}\pm 2.420\mathrm{e}^{-3}}$ $\bm{4.554\mathrm{e}^{-2}\pm 1.119\mathrm{e}^{-3}}$ $100$
Figure 10: Stonehenge Scene with $\delta_{R}=20^{\circ}$ and $\delta_{t}=0.1\,\mathrm{m}$—While ICP and Colored-ICP diverge over time, Splat-Loc-SIFT and Splat-Loc-Glue do not diverge, producing high-accuracy estimates.
In addition, we examine the performance of the pose estimation algorithmsin problems with a larger error in the initial estimate of the pose, with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$. We present the performance of each algorithm on each metric in Table IX, where we note that ICP and Colored-ICP do not provide accurate estimate of the robot’s pose. Moreover, the pose estimation errors achieved by ICP and Colored-ICP have a significant variance. In contrast, Splat-Loc-SIFT and Splat-Loc-Glue yield pose estimates of high accuracy with average rotation and translation errors less than $0.5$ deg. and 55mm, respectively, similar to the setting with low errors in the initial pose estimate. However, Splat-Loc-SIFT achieves a lower success rate, compared to Splat-Loc-Glue, which achieves a perfect success rate. In Figure 11, we show the estimation errors for a single run, highlighting the high variance in the errors associated with the pose estimates computed Colored-ICP. ICP fails to yield a pose estimate of relatively high accuracy for the duration of the problem.
@@
TABLE IX: Comparison of pose estimation algorithms in the Stonehenge scene with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$.
@@
Algorithm Rotation (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$) ICP [4] $1.303\mathrm{e}^{2}\pm 2.254\mathrm{e}^{1}$ $3.697\mathrm{e}^{0}\pm 5.533\mathrm{e}^{0}$ $1.217\mathrm{e}^{-1}\pm 1.527\mathrm{e}^{-1}$ $100$ Colored-ICP [34] $9.489\mathrm{e}^{1}\pm 5.128\mathrm{e}^{1}$ $5.737\mathrm{e}^{-1}\pm 2.880\mathrm{e}^{-1}$ $4.879\mathrm{e}^{-1}\pm 1.038\mathrm{e}^{-1}$ 2020 Splat-Loc-SIFT (ours) $\bm{2.164\mathrm{e}^{-1}\pm 3.682\mathrm{e}^{-2}}$ $3.335\mathrm{e}^{-3}\pm 5.628\mathrm{e}^{-4}$ $1.384\mathrm{e}^{-1}\pm 3.145\mathrm{e}^{-3}$ $70$ Splat-Loc-Glue (ours) $2.193\mathrm{e}^{-1}\pm 2.024\mathrm{e}^{-1}$ $\bm{3.141\mathrm{e}^{-3}\pm 2.099\mathrm{e}^{-3}}$ $\bm{4.507\mathrm{e}^{-2}\pm 6.105\mathrm{e}^{-4}}$ $100$
Figure 11: Stonehenge Scenewith $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$Splat-Loc-SIFT and Splat-Loc-Glue produce pose estimates of the highest accuracy, compared to ICP and Colored-ICP. Further, Splat-Loc-Glue requires the least computation time per frame.

### v2

@@ C Discussion of Safety
D Splat-Plan MPC
Since the polytope and A* computation procedures are relatively fast, while computation of the B-spline is relatively slow, we can compute the B-spline for some small time step in the future, in a receding-horizon manner. In our Splat-Plan MPC, we do this by fitting a B-spline through the current and next polytope. This still guarantees safety and feasibility since the entire constraint set is separable based on time, at the expense of the optimality of the solution.
TABLE IV: Comparison of Local Trajectory planning algorithm computation times.
@@ D Splat-Plan MPC
To push the limits of generalizability of the polytopes, we implemented a simple line following planner based on the polytopes. Specifically, in this baseline, we query a polytope at the robot’s current position, and another polytope at a point some small distance away ($0.02$ in our evaluations) on the line between the current position and the goal position, then solve for a B-spline. Such a method did not yield connected polytopes for $18\%$ of configurations, and got stuck (ran for over $200$ iterations) for $24\%$ of configurations. Note that while the robot was stuck, the Bézier spline was still guaranteed safe (Table IV).
Appendix G Recursive Pose Estimation
We present additional results demonstrating the performance of Splat-Loc compared to other algorithms. Table V shows our performance metrics in the Statues scene with with $\delta_{R}=20^{\circ}$} and $\delta_{t}=0.1\,\mathrm{m}$. First, we see that only the Colored-ICP algorithm fails to achieve a $100\%$ success rate. Second, our algorithms Splat-Loc-Glue and Splat-Loc-SIFT yield higher-accuracy and more consistent solutions, with mean rotation and translation errors less than 0.1 deg and 7 mm, respectively, and lower standard deviations. In these tests, Splat-Loc-Glue performs the best on each metric with a computation time of about 35 ms.
@@ Appendix G Recursive Pose Estimation
TABLE V: Comparison of pose estimation algorithms in the Statues scene with $\delta_{R}=20^{\circ}$ and $\delta_{t}=0.1\,\mathrm{m}$
@@
Algorithm Rotation Error (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$)
ICP [54] $8.42\mathrm{e}^{1}\pm 5.58\mathrm{e}^{1}$ $3.83\mathrm{e}^{0}\pm 8.51\mathrm{e}^{0}$ $8.60\mathrm{e}^{-2}\pm 4.40\mathrm{e}^{-2}$ $100$
Colored-ICP [50] $3.02\mathrm{e}^{1}\pm 4.58\mathrm{e}^{1}$ $3.46\mathrm{e}^{-1}\pm 5.11\mathrm{e}^{-1}$ $6.27\mathrm{e}^{-2}\pm 2.72\mathrm{e}^{-2}$ $80$
Splat-Loc-SIFT (ours) $8.32\mathrm{e}^{-2}\pm 9.37\mathrm{e}^{-3}$ $5.72\mathrm{e}^{-3}\pm 7.62\mathrm{e}^{-4}$ $6.56\mathrm{e}^{-2}\pm 7.16\mathrm{e}^{-4}$ $100$
Splat-Loc-Glue (ours) $\bm{5.73\mathrm{e}^{-2}\pm 4.82\mathrm{e}^{-3}}$ $\bm{3.13\mathrm{e}^{-3}\pm 3.24\mathrm{e}^{-4}}$ $\bm{3.42\mathrm{e}^{-2}\pm 2.00\mathrm{e}^{-3}}$ $100$
In Table VI, we present results for the Statues scene with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$. Note that the success rates of Colored-ICP and Splat-Loc-SIFT decrease sharply, suggesting that these methods require a good initial estimate of the pose of the robot. However, when they do succeed, these algorithms achieve low rotation and translation errors, with Splat-Loc-SIFT achieving a much-higher accuracy compared to Colored-ICP. In contrast, the Splat-Loc-Glue algorithm achieves a perfect success rate, along with the lowest rotation and translation errors and computation time. The results suggest the greater robustness of Splat-Loc-Glue compared to Splat-Loc-SIFT.
@@
TABLE VI: Comparison of pose estimation algorithms in the Statues scene with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$
@@
Algorithm Rotation Error (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$)
ICP [54] $9.90\mathrm{e}^{1}\pm 3.06\mathrm{e}^{1}$ $1.41\mathrm{e}^{0}\pm 2.30\mathrm{e}^{-1}$ $4.36\mathrm{e}^{-2}\pm 5.08\mathrm{e}^{-2}$ $100$
Colored-ICP [50] $2.42\mathrm{e}^{0}\pm 1.61\mathrm{e}^{0}$ $2.19\mathrm{e}^{-1}\pm 2.02\mathrm{e}^{-1}$ $7.03\mathrm{e}^{-2}\pm 2.44\mathrm{e}^{-2}$ 2020
Splat-Loc-SIFT (ours) $7.85\mathrm{e}^{-2}\pm 5.96\mathrm{e}^{-3}$ $6.28\mathrm{e}^{-3}\pm 8.40\mathrm{e}^{-4}$ $7.32\mathrm{e}^{-2}\pm 1.50\mathrm{e}^{-3}$ $30$
Splat-Loc-Glue (ours) $\bm{6.14\mathrm{e}^{-2}\pm 6.00\mathrm{e}^{-3}}$ $\bm{3.68\mathrm{e}^{-3}\pm 6.54\mathrm{e}^{-4}}$ $\bm{4.28\mathrm{e}^{-2}\pm 1.50\mathrm{e}^{-3}}$ $100$
In addition, we examine the performance of the pose estimation algorithms in problems with a larger error in the initial estimate of the pose, with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$ in the synthetic Stonehenge scene. We present the performance of each algorithm on each metric in Table VII, where we note that ICP and Colored-ICP do not provide accurate estimate of the robot’s pose. Moreover, the pose estimation errors achieved by ICP and Colored-ICP have a significant variance. In contrast, Splat-Loc-SIFT and Splat-Loc-Glue yield pose estimates of high accuracy with average rotation and translation errors less than $0.5$ deg. and 55mm, respectively, similar to the setting with low errors in the initial pose estimate. However, Splat-Loc-SIFT achieves a lower success rate, compared to Splat-Loc-Glue, which achieves a perfect success rate.
TABLE VII: Comparison of pose estimation algorithms in the Stonehenge scene with $\delta_{R}=30^{\circ}$ and $\delta_{t}=0.5\,\mathrm{m}$.
Algorithm Rotation Error (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$)
ICP [54] $1.31\mathrm{e}^{2}\pm 2.26\mathrm{e}^{1}$ $3.70\mathrm{e}^{0}\pm 5.54\mathrm{e}^{0}$ $1.22\mathrm{e}^{-1}\pm 1.53\mathrm{e}^{-1}$ $100$
Colored-ICP [50] $9.49\mathrm{e}^{1}\pm 5.13\mathrm{e}^{1}$ $5.74\mathrm{e}^{-1}\pm 2.88\mathrm{e}^{-1}$ $4.88\mathrm{e}^{-1}\pm 1.04\mathrm{e}^{-1}$ 2020
Splat-Loc-SIFT (ours) $\bm{2.17\mathrm{e}^{-1}\pm 3.69\mathrm{e}^{-2}}$ $3.34\mathrm{e}^{-3}\pm 5.63\mathrm{e}^{-4}$ $1.39\mathrm{e}^{-1}\pm 3.15\mathrm{e}^{-3}$ $70$
Splat-Loc-Glue (ours) $2.20\mathrm{e}^{-1}\pm 2.03\mathrm{e}^{-1}$ $\bm{3.15\mathrm{e}^{-3}\pm 2.10\mathrm{e}^{-3}}$ $\bm{4.51\mathrm{e}^{-2}\pm 6.11\mathrm{e}^{-4}}$ $100$
Appendix H Global Pose Initialization of Splat-Loc
We evaluate the global point-cloud alignment initialization procedure from Section V-E when utilized by Splat-Loc-Glue, in the real-world scene Statues and synthetic scene Stonehenge, with the results in Table VIII. We see that the success rate is 80% in both scenes, and it runs at approximately 20-30 Hz. When it does succeed, the pose errors are on the order of 0.3 deg and 1 cm. Although these estimates are relatively good, we note that these are less accurate than the recursive pose estimates discussed above.
TABLE VIII: Performance of the pose initialization module in Splat-Loc
Scene Rotation Error (deg.) Translation Error (m) Computation Time (secs.) Success Rate ($\%$)
Statues $3.69\mathrm{e}^{-1}\pm 7.21\mathrm{e}^{-1}$ $1.34\mathrm{e}^{-2}\pm 2.33\mathrm{e}^{-2}$ $3.39\mathrm{e}^{-2}\pm 5.70\mathrm{e}^{-4}$ $80$
Stonehenge $2.14\mathrm{e}^{-1}\pm 2.07\mathrm{e}^{-1}$ $3.03\mathrm{e}^{-3}\pm 2.28\mathrm{e}^{-3}$ $4.58\mathrm{e}^{-2}\pm 1.70\mathrm{e}^{-3}$ $80$